OpenStudy (jkbo):

Please help me A.3m/5 + a = b − 2m/5 5/3a − b B.3/5a − b C.3/5b − a D.m = b − a

OpenStudy (anonymous):

give me the equation

OpenStudy (jkbo):

Sorry I wrote it wrong @samsterz 3m/5 + a = b − 2m/5 A.5/3a − b B.3/5a − b C.3/5b − a D.m = b − a